# LLMs instead of Human Judges? A Large Scale Empirical Study across 20 NLP Evaluation Tasks

### v2

@@ A Large Scale Empirical Study across 20 NLP Evaluation Tasks
Anna Bavaresco1, Raffaella Bernardi2, Leonardo Bertolazzi2, Desmond Elliott3, Raquel Fernández1, Albert Gatt4, Esam Ghaleb5, Mario Giulianelli6, Michael Hanna1, Alexander Koller7, André F. T. Martins8, Philipp Mondorf9, Vera Neplenbroek1, Sandro Pezzelle1, Barbara Plank9, David Schlangen10, Alessandro Suglia11, Aditya K Surikuchi1, Ece Takmaz4, Alberto Testoni1 1University of Amsterdam, 2University of Trento, 3University of Copenhagen, 4Utrecht University, 5Max Planck Institute for Psycholinguistics, 6ETH Zürich, 7Saarland University, 8Universidade de Lisboa & Unbabel, 9LMU Munich & MCML, 10University of Potsdam, 11Heriot-Watt University
@@ 1 Introduction
For many natural language processing (NLP) tasks, the most informative evaluation is to ask humans to judge the model output. Such judgments are traditionally collected in lab experiments or through crowdsourcing, with either expert or non-expert annotators, as illustrated in Fig. 1. Recently, there has been a trend towards replacing human judgments with automatic assessments obtained via large language models (LLMs) (Chiang and Lee, 2023; Wang et al., 2023a; Liu et al., 2023; Li et al., 2024; Zheng et al., 2024, inter alia). For example, the LLM could be instructed to rate a response generated by a dialogue system for its perceived plausibility on a scale from 1 to 5. This drastically reduces the evaluation effort and is claimed to yield more reliable results across multiple evaluation rounds Landwehr et al. (2023); Jiang et al. (2023b); Reiter (2024); Dubois et al. (2024).
@@
Previous studies offer mixed evidence regarding the reliability of LLM evaluators. Some research concludes that they are effective, correlating well with human judgments (Liu et al., 2023; Zheng et al., 2024; Chen et al., 2023; Verga et al., 2024; Törnberg, 2023; Huang et al., 2024; Naismith et al., 2023; Gilardi et al., 2023; Kocmi and Federmann, 2023b), albeit with some caveats (Wang et al., 2023a; Wu and Aji, 2023; Hada et al., 2024; Pavlovic and Poesio, 2024). In some cases, LLM evaluators can also provide pairwise preference judgments Kim et al. (2024); Liusie et al. (2024); Liu et al. (2024a); Park et al. (2024); Tan et al. (2024) or fine-grained evaluation beyond a single score, such as error spans (Fernandes et al., 2023; Kocmi and Federmann, 2023a). In contrast, some studies highlight substantial biases in LLMs’ behaviour as evaluators, both as compared against human judgments (Koo et al., 2023; Zeng et al., 2024; Baris Schlicht et al., 2024) and through intrinsic analyses Wang et al. (2023b); Liu et al. (2024b); Stureborg et al. (2024). These discrepancies likely stem from the limitations of this previous work, which typically relies on a few datasets and models, often restricted to closed-source proprietary models.
In this paper, we examine how well current LLMs can approximate human evaluators on a large scale. We prompt 11 among the most recent open-weight and proprietary LLMs to generate judgments on 20 datasets with human annotations on a wide range of quality dimensions, prompt styles, and tasks. Our evaluation goes beyond existing work by including a wide variety of datasets that differ in the type of task (e.g., translation, dialogue generation, etc.), the property being judged (e.g., coherence, fluency, etc.), the type of judgments (categorical or graded), and the expertise of human annotators (experts or non-experts). We provide Judge-Bench, a benchmark which includes upon release a total of over 70,000 test instances with associated human judgments with an extensible codebase. [^1] https://github.com/dmg-illc/JUDGE-BENCH Our results indicate that LLMs align well with human judgments on certain tasks, like instruction following. However, their performance is inconsistent across and within annotation tasks. Elicitation methods like Chain-of-Thought prompting do not reliably improve agreement, in line with recent findings Sprague et al. (2024). Some proprietary models—in particular, GPT-4o—align better to humans, but there is a rather small gap with large open-source models, holding promise for the reproducibility of future evaluation efforts. Altogether, at the current stage of LLM development, we recommend validating LLM judges against task-specific human annotations before deploying them for any particular task.
@@ 2 Construction of Judge-Bench
One key feature that differs across the datasets included in Judge-Bench is the source of the data being evaluated, i.e., whether the items to be judged are generated by a model or produced by humans (Fig. 1). For model-generated items, the goal is to evaluate an NLP system. This includes both classic tasks such as machine translation or dialogue response generation, as well as less standard tasks for which automation has recently become an option thanks to LLMs, such as the generation of plans or logical arguments. For human-generated items, the goal is to assess properties of interest such as grammaticality or toxicity. This distinction allows us to understand whether LLMs have a positive bias towards machine-generated outputs—a tendency reported in prior work Xu et al. (2024).
@@
Table 1: Scores per dataset for the models with $\geq$98valid response rates (results for all models in Tab. 5, App. F): Cohen’s kappa for categorical annotations and Spearman’s correlation for graded annotations. Boldface marks best model performance per dataset. Datasets with both categorical and graded annotations appear twice. Datasets in blue concern human-generated language, while those in red concern model-generated text. ‘$\sigma$’ denotes the standard deviation of the scores across models per dataset (averaged over properties if more than one is judged per dataset). Upper-bound estimates (UB) indicate the agreement between individual and aggregated human judgments.
Our study focuses on English datasets or language pairs which include English as one of the languages. We keep track of whether the original annotation guidelines are available and whether the annotations are provided by expert or non-expert annotators. We retain all available individual annotations. Dataset information is summarised in Tab. 2, App. A. All 20 datasets are formatted following a precise data schema to facilitate the integration of additional datasets. This makes Judge-Bench easily extensible.
@@ 3 Model Selection and Experiment Design
We select representative proprietary and open-weight models of various sizes that show high performance across several tasks on the Open LLM and Chatbot Arena Leaderboards Chiang et al. (2024): GPT-4o OpenAI (2024), LLaMA-3.1 (8B and 70B;  AI@Meta 2024), Gemini-1.5 Reid et al. (2024), Mixtral (8x7B and 8x22B; Jiang et al. 2024), Command R and Command R+ Cohere and Cohere for AI (2024a, b), OLMo Groeneveld et al. (2024), Starling-7B Zhu et al. (2023), and Mistral Jiang et al. (2023a). See App. C for inference procedure details.
@@
Since most datasets include the original instructions used to gather human judgments, we use these instructions directly as prompts for the model, with additional guidelines to constrain the models’ output and minimise verbosity: ‘Answer with one of {}. Do not explain your answer.’ When the original instruction for collecting human judgments is unavailable, we create a prompt based on relevant information from the original paper, such as the task description and the definitions of the evaluation metrics. We also experimented with alternative prompting strategies, including Chain-of-Thought, few-shot and system prompts, and prompt paraphrases. We do not observe systematic improvements. See App. F for full details and results. All prompts are provided in the codebase.
@@
Models do not always respond to the prompts as requested (e.g., they may refuse to answer if they perceive the prompt as sensitive). We therefore use the following evaluation protocol: ($i$) To obtain the same number of judgments across models for a given dataset, we replace invalid LLM responses with judgments randomly sampled from the relevant set of categorical or graded annotations. Fig. 4 in App. D shows the rate of valid responses per model. ($ii$) For graded annotations, we compute Spearman’s correlation ($\rho$) between model and human judgments; for categorical annotations, we compute Cohen’s $\kappa$. ($iii$) When multiple individual human judgments are available, we estimate an upper bound by computing the average Spearman’s $\rho$ or Cohen’s $\kappa$ between bootstrapped single-rater responses and the aggregated responses across raters.222 More details on the upper bound calculation are in App. B. Tab. 3 (App. A) reports Krippendorff’s $\alpha$. Datasets containing multiple human judgments are marked in Tab. 2 (App. A).
@@ 4 Results
Scores vary substantially across models. For any given model, they vary both across datasets and properties being judged. Tab. 1 presents detailed results for the 6 models that exhibit the largest rate of valid responses ($\geq$98%). GPT-4o ranks first across several evaluation scenarios, but the Llama-3.1-70B and Mixtral-8x22B open models are relatively close and outperform GPT-4o on some assessment types, such as categorical sentence acceptability (CoLa) and graded summary quality (SummEval). Overall, the high degree of variability is not fully accounted for by the inherent difficulty of the annotation tasks as reflected in the human upper bound. Moreover, except for a few datasets (e.g., QAGS, Recipe-generation, and NewsRoom), model scores remain notably below the upper bound.
Among the property types with the lowest human-model alignment are toxicity and safety (in particular on DICES and Medical-safety), where model scores can be even negative and valid response rates particularly low (see Fig. 5 in App. D). This is due in part to the guardrails associated with these tasks Weidinger et al. (2023). We find that, especially in the medical domain, many models tend to provide explanations instead of producing a judgment (see App. E).Despite the high variability across models and datasets, we observe several notable trends. For graded annotations (Fig. 2), all models achieve higher correlations with annotations by non-expert human judges compared to expert annotators, echoing recent findings by Aguda et al. (2024).
Figure 3 shows correlation results across different datasets for the subset of properties that exclusively have graded judgments. The proprietary models GPT-4o and Gemini-1.5 exhibit the highest scores when evaluating acceptability and verbosity, while the two Mixtral open models show the strongest correlations for coherence and consistency. Overall, no single model demonstrates a clear superiority over others across all categories; instead, different quality dimensions are better assessed by different models.
Finally, all models achieve better alignment with human judgments when evaluating human language than when assessing machine-generated text, both for categorical and graded annotations (see Fig. 6 in App. F). This emphasises the need for caution when using LLMs to automatically evaluate the output of NLP systems.
@@ 5 Conclusions
In response to current trends in evaluation, in this paper we conducted a large-scale study of the correlation between human and LLM judgments across 20 datasets, considering factors such as the properties being assessed, the expertise level of the human judges, and whether the data is model- or human-generated. On some tasks, such as instruction following and the generation of mathematical reasoning traces, models can be reliably used as evaluators. Overall, however, models’ agreement with human judgments varies widely across datasets, evaluated properties, and data sources; and elicitation strategies such as Chain-of-Thought prompting do not consistently improve agreement levels, in line with recent findings Sprague et al. (2024). We recommend validation and calibration of LLMs against task-specific human judgments prior to their deployment as evaluators. To facilitate this process, we release Judge-Bench, a benchmark that enables systematic evaluation across a diverse range of tasks and is easily extensible to include any new task of interest.
@@
As pointed out by one of the reviewers, correlation with human judges may not be the most appropriate way to validate LLM evaluators. Indeed, if the responses of an LLM were found to contain some harmful bias that does not affect the overall correlation or to be systematically aligned with the beliefs of one specific group (without taking into account other perspectives), this would arguably not be a good reason to conclude that LLMs are good evaluators. However, we believe that there are tasks where it is still useful and informative to compare LLM judgments against human ones, especially if human annotations come from experts. The reviewer also highlights the potential dangers of reusing pre-existing tasks and datasets without verifying their quality or how well they reflect actual downstream tasks. While we did our best to select a set of tasks that would be representative and meaningful for the NLP community, we acknowledge that there are potential shortcomings (such as data leakage) in using pre-existing tasks and datasets without revalidating them.
In contrast to approaches that use LLMs for pairwise preference evaluation, e.g., PairEval (Park et al., 2024) or JudgeBench Tan et al. (2024), this paper focuses on evaluating the performance of LLMs on generating judgements for categorical or graded responses. We leave extending Judge-Bench to include pairwise preference evaluation and other recent evaluation methods like Prometheus 2 (Kim et al., 2024) to future work.

### v3

@@ A Large Scale Empirical Study across 20 NLP Evaluation Tasks
Anna Bavaresco1, Raffaella Bernardi2, Leonardo Bertolazzi2, Desmond Elliott3, Raquel Fernández1, Albert Gatt4, Esam Ghaleb5, Mario Giulianelli6, Michael Hanna1, Alexander Koller7, André F. T. Martins8, Philipp Mondorf9, Vera Neplenbroek1, Sandro Pezzelle1, Barbara Plank9, David Schlangen10, Alessandro Suglia11, Aditya K Surikuchi1, Ece Takmaz4, Alberto Testoni12 1University of Amsterdam, 2University of Trento, 3University of Copenhagen, 4Utrecht University, 5Max Planck Institute for Psycholinguistics, 6ETH Zürich, 7Saarland University, 8Universidade de Lisboa & Unbabel, 9LMU Munich & MCML, 10University of Potsdam, 11Heriot-Watt University 12Amsterdam UMC judgebench@googlegroups.comthanks: Authors listed in alphabetical order.
@@ 1 Introduction
For many natural language processing (NLP) tasks, the most informative evaluation is to ask humans to judge the model output. Such judgments are traditionally collected in lab experiments or through crowdsourcing, with either expert or non-expert annotators, as illustrated in Fig. 1. Recently, there has been a trend towards replacing human judgments with automatic assessments obtained via large language models (LLMs; Chiang and Lee, 2023; Wang et al., 2023; Liu et al., 2023; Li et al., 2024; Zheng et al., 2024, inter alia). For example, the LLM could be instructed to rate a response generated by a dialogue system for its perceived plausibility, on a scale from 1 to 5. This drastically reduces the evaluation effort and is claimed to yield more reliable results across multiple evaluation rounds Landwehr et al. (2023); Jiang et al. (2023b); Reiter (2024); Dubois et al. (2024).
@@
Previous studies offer mixed evidence regarding the reliability of LLM evaluators. Some research concludes that they are effective, correlating well with human judgments (Liu et al., 2023; Zheng et al., 2024; Chen et al., 2023; Verga et al., 2024; Törnberg, 2023; Huang et al., 2024; Naismith et al., 2023; Gilardi et al., 2023; Kocmi and Federmann, 2023b), albeit with some caveats (Wang et al., 2023; Wu and Aji, 2025; Hada et al., 2024; Pavlovic and Poesio, 2024). In some cases, LLM evaluators can also provide pairwise preference judgments Kim et al. (2024); Liusie et al. (2024); Liu et al. (2024a); Park et al. (2024); Tan et al. (2025), or fine-grained evaluation beyond a single score, such as error spans (Fernandes et al., 2023; Kocmi and Federmann, 2023a). In contrast, some studies highlight substantial biases in LLMs’ behaviour as evaluators, both as compared against human judgments (Koo et al., 2024; Zeng et al., 2024; Baris Schlicht et al., 2024) and through intrinsic analyses Wang et al. (2024); Liu et al. (2024b); Stureborg et al. (2024). These discrepancies likely stem from the limitations of this previous work, which typically relies on a few datasets and models, often restricted to closed-source proprietary models. The observation of such limitations has motivated recent work to develop finetuning methods for LLM judges designed to overcome certain biases Zhu et al. (2025).
In this paper, we examine how well current LLMs can approximate human evaluators on a large scale. We prompt 11 among the most recent open-weight and proprietary LLMs to generate judgments on 20 datasets with human annotations on a wide range of quality dimensions, prompt styles, and tasks. Our evaluation goes beyond existing work by including a wide variety of datasets that differ in the type of task (e.g., translation, dialogue generation, etc.), the property being judged (e.g., coherence, fluency, etc.), the type of judgments (categorical or graded), and the expertise of human annotators (experts or non-experts). We provide Judge-Bench, a benchmark which includes, upon release, a total of over 70,000 test instances with associated human judgments with an extensible codebase.
Our results indicate that LLMs align well with human judgments on certain tasks, like instruction following. However, their performance is inconsistent across and within annotation tasks. Elicitation methods like Chain-of-Thought prompting Wei et al. (2022) do not reliably improve agreement, in line with recent findings Sprague et al. (2025). Some proprietary models—in particular, GPT-4o—align better to humans, but there is a rather small gap with large open-source models, holding promise for the reproducibility of future evaluation efforts. Altogether, at the current stage of LLM development, we recommend validating LLM judges against task-specific human annotations before deploying them for any particular task.
@@ 2 Construction of Judge-Bench
One key feature that differs across the datasets included in Judge-Bench is the source of the data being evaluated, i.e., whether the items to be judged are generated by a model or produced by humans, as illustrated in Figure 1.
For model-generated items, the goal is to evaluate an NLP system. This includes both classic tasks such as machine translation or dialogue response generation, as well as less standard tasks for which automation has recently become an option thanks to LLMs, such as the generation of plans or logical arguments. For human-generated items, the goal is to assess properties of interest such as grammaticality or toxicity. This distinction allows us to understand whether LLMs have a positive bias towards machine-generated outputs—a tendency reported in prior work Xu et al. (2024).
@@
Table 1: Scores per dataset for the models with $\geq$98valid response rates (results for all models in Tab. 6, App. H): Cohen’s kappa for categorical annotations and Spearman’s correlation for graded annotations. Boldface marks best model performance per dataset. Spearman’s correlations are generally significant ($p<0.05$), with the exception of the Persona Chat and Topical Chat datasets (see Tab. 6 in Appendix H for more details). Datasets with both categorical and graded annotations appear twice. Datasets in blue concern human-generated language, while those in red concern model-generated text. ‘$\sigma$’ denotes the standard deviation of the scores across models per dataset (averaged over properties if more than one is judged per dataset). Upper-bound estimates (UB) indicate the agreement between individual and aggregated human judgments.
Our study focuses on English datasets or language pairs which include English as one of the languages. We keep track of whether the original annotation guidelines are available and whether the annotations are provided by experts or non-experts. We retain all available individual annotations. Dataset information is summarised in Table 2, Appendix A. All 20 datasets are formatted following a precise data schema to facilitate the integration of additional datasets. This makes Judge-Bench easily extensible. We provide more details about the data schema in Appendix B.
@@ 3 Model Selection and Experiment Design
We select representative proprietary and open-weight models of various sizes that show high performance across several tasks on the Open LLM and Chatbot Arena Leaderboards Chiang et al. (2024): GPT-4o OpenAI (2024), LLaMA-3.1 (8B and 70B;  AI@Meta 2024), Gemini-1.5 Reid et al. (2024), Mixtral (8x7B and 8x22B; Jiang et al. 2024), Command R and Command R+ Cohere and Cohere for AI (2024a, b), OLMo Groeneveld et al. (2024), Starling-7B Zhu et al. (2024), and Mistral Jiang et al. (2023a). See Appendix E for inference procedure details.
@@
Since most datasets include the original instructions used to gather human judgments, we use these instructions directly as prompts for the model, with additional guidelines to constrain the models’ output and minimise verbosity: ‘Answer with one of {}. Do not explain your answer.’ When the original instruction for collecting human judgments is unavailable, we create a prompt based on relevant information from the original paper, such as the task description and the definitions of the evaluation metrics. We also experiment with alternative prompting strategies, including Chain-of-Thought, few-shot prompts, and prompt paraphrases. However, none of these strategies leads to systematic improvements. See Appendix H for full details and results. All prompts are provided in the codebase.
@@
Models do not always respond to the prompts as requested (e.g., they may refuse to answer if they perceive the prompt as sensitive). We therefore use the following evaluation protocol:
  • To obtain the same number of judgments across models for a given dataset, we replace invalid LLM responses with judgments randomly sampled from the relevant set of categorical or graded annotations. Figure 5 in Appendix F shows the rate of valid responses per model.
  • Graded annotations, such as in WMT 2020 (Freitag et al., 2021), assess language quality on a continuous scale (e.g., a score from 0 to 100, or Likert-scale ratings), capturing varying degrees of fluency, adequacy, or overall translation quality; whereas categorical annotations, like those in CoLa (Warstadt et al., 2019), involve binary judgments (e.g., grammatically acceptable or not). For the former, we compute Spearman’s correlation ($\rho$) between model and human judgments; for the latter, we compute Cohen’s $\kappa$.
  • When multiple individual human judgments are available (typically three, see Table 2 in Appendix A), we estimate an upper bound by computing the average Spearman’s $\rho$ or Cohen’s $\kappa$ between bootstrapped single-rater responses and the aggregated responses across raters. Appendix C provides details on the upper bounds.
@@ 4 Results
Scores vary substantially across models. For any given model, they vary both across datasets and properties being judged. Table 1 presents detailed results for the 6 models that exhibit the largest rate of valid responses ($\geq$98%). GPT-4o ranks first across several evaluation scenarios, but the Llama-3.1-70B and Mixtral-8x22B open models are relatively close and outperform GPT-4o on some assessment types, such as categorical sentence acceptability (CoLa) and graded summary quality (SummEval). Overall, the high degree of variability is not fully accounted for by the inherent difficulty of the annotation tasks, as reflected in the human upper bound. Moreover, except for a few datasets (e.g., QAGS, Recipe-generation, and NewsRoom), model scores remain notably below the upper bound.
Among the property types with the lowest human-model alignment are toxicity and safety (in particular on DICES and Medical-safety), where model scores can be even negative and valid response rates particularly low (see Fig. 6 in Appendix F). This is due in part to the guardrails associated with these tasks Weidinger et al. (2023). We find that, especially in the medical domain, many models tend to provide explanations instead of producing a judgment (see Appendix G).
Despite the high variability across models and datasets, we observe several notable trends. For graded annotations (Fig. 2), all models achieve higher correlations with annotations by non-expert human judges compared to expert annotators, echoing recent findings by Aguda et al. (2024). One possible explanation is that non-experts might rely on surface-level features, which could align more closely with the patterns LLMs are most attuned to, while experts apply stricter, domain-specific criteria. This remains speculative and calls for further investigation.
Figure 3 shows correlation results across different datasets for the subset of properties that exclusively have graded judgments. When applicable, we average results across datasets including annotations for the same property. We provide more details about these properties in Appendix D. The proprietary models GPT-4o and Gemini-1.5 exhibit the highest scores when evaluating acceptability and verbosity, while the two Mixtral open models show the strongest correlations for coherence and consistency. Correlation with the engagingness property remains consistently low across all models. Overall, no single model demonstrates a clear superiority over others across all properties; instead, different quality dimensions are better assessed by different models. This calls into question the widespread practice of using a single model – typically a proprietary one like those from the GPT family—to evaluate a diverse range of linguistic properties.
Finally, as shown in Figure 4, all models achieve better alignment with human judgments when evaluating human language than when assessing machine-generated text, both for categorical and graded annotations. This result aligns with the findings by Xu et al. (2024), suggesting that LLMs display a bias towards their own generation. More broadly, this trend calls for caution when using LLMs to automatically evaluate the output of NLP systems.
@@ 5 Conclusions
In response to current trends in evaluation, in this paper we conducted a large-scale study of the correlation between human and LLM judgments across 20 datasets, considering factors such as the properties being assessed, the expertise level of the human judges, and whether the data is model- or human-generated. On some tasks, such as instruction following and the generation of mathematical reasoning traces, models can be reliably used as evaluators. Overall, however, models’ agreement with human judgments varies widely across datasets, evaluated properties, and data sources; and depends on the level of expertise of human judges. Furthermore, elicitation strategies such as Chain-of-Thought prompting do not consistently improve agreement levels, in line with recent findings Sprague et al. (2025). We recommend validation and calibration of LLMs against task-specific human judgments prior to their deployment as evaluators. To facilitate this process, we release Judge-Bench, a benchmark that enables systematic evaluation across a diverse range of tasks and is easily extensible to include any new task of interest.
@@
One limitation of the experimental design of our work is that correlation with human judges may not be the most appropriate way to validate LLM evaluators. Indeed, there may be domains where human annotators and LLM evaluators appear aligned simply because they are affected by similar biases. Therefore, depending on the task at hand, it may be necessary to validate the reliability of human annotators as well.
Another limitation concerns the use of existing tasks and datasets without reassessing their quality or representativeness of actual downstream tasks. While we did our best to select a wide set of tasks meaningful to the NLP community, we acknowledge that these tasks could not be equally meaningful for end-users, and that employing existing datasets could arguably lead to potential risks and shortcomings, such as data leakage.
In contrast to approaches that use LLMs for pairwise preference evaluation, e.g., PairEval (Park et al., 2024) or JudgeBench Tan et al. (2025), [^1] Some time after an early version of this paper became available as a pre-print, accompanied by our Judge-Bench code, the independent work by Tan et al. (2025) appeared, which describes a benchmark that the authors named JudgeBench. This name clash is unfortunate, but since in the meantime our paper has seen some uptake, we have decided against trying to resolve it. this paper focuses on evaluating the performance of LLMs on generating judgements for categorical and graded responses. We leave the extension of Judge-Bench to include pairwise preference evaluation and other recent evaluation methods, such as Prometheus 2 (Kim et al., 2024), for future work.
@@
Acknowledgements
This work emerged from discussions at a workshop organised by Raquel Fernández and Sandro Pezzelle at MFO, the Oberwolfach Research Institute for Mathematics in the German Black Forest, on behalf of the ELLIS NLP programme. The event was funded by the state of Baden-Württemberg (Germany) and organised in collaboration with the ELLIS Institute Tübingen and the Max Planck Institute for Intelligent Systems. We furthermore acknowledge our funders. In particular, AB, RF, and AT were supported by the European Research Council (ERC Consolidator Grant DREAM 819455 to RF), as well as BP (ERC Consolidator Grant DIALECT 101043235 to BP). DE was supported by a research grant (VIL53122) from VILLUM FONDEN. EG was supported by the Dutch Research Council (Gravitation grant 024.001.006 to the Language in Interaction consortium). MG was supported by an ETH Zurich Postdoctoral Fellowship. MH was supported in part by an OpenAI Superalignment Fellowship. AM was supported by the European Research Council (DECOLLAGE, ERC-2022-CoG 101088763) and by Fundação para a Ciência e Tecnologia through contract UIDB/50008/2020. We acknowledge ISCRA for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy).